OpenStudy (anonymous):

Use the quadratic formula to solve the following quadratic equations: 1. 3x = 4 - x^2 2. 2x^2 - 6 = 3x

OpenStudy (unklerhaukus):

first get all the terms on the left hand side of the equation \[ax^2+bx+c=0\] then apply the quadratic formula \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]